# Rethinking Gating Mechanism in Sparse MoE: Handling Arbitrary Modality Inputs with Confidence-Guided Gate

### v1

@@ Abstract
Effectively managing missing modalities is a fundamental challenge in real-world multimodal learning scenarios, where data incompleteness often results from systematic collection errors or sensor failures. Sparse Mixture-of-Experts (SMoE) architectures have the potential to naturally handle multimodal data, with individual experts specializing in different modalities. However, existing SMoE approach often lacks proper ability to handle missing modality, leading to performance degradation and poor generalization in real-world applications. We propose Conf-SMoE to introduce a two-stage imputation module to handle the missing modality problem for the SMoE architecture and reveal the insight of expert collapse from theoretical analysis with strong empirical evidence. Inspired by our theoretical analysis, Conf-SMoE propose a novel expert gating mechanism by detaching the softmax routing score to task confidence score w.r.t ground truth. This naturally relieves expert collapse without introducing additional load balance loss function. We show that the insights of expert collapse aligns with other gating mechanism such as Gaussian and Laplacian gate. We also evaluate the proposed method on four different real world dataset with three different experiment settings to conduct comprehensive the analysis of Conf-SMoE on modality fusion and resistance to missing modality. We release our official implementation in: https://github.com/IcurasLW/Official-Repository-of-ConfSMoE.git
@@ 1 Introduction
As modern applications increasingly involve data from multiple heterogeneous sources, multimodal learning has emerged as a critical paradigm to unlock complementary insights and enhance decision-making in complex tasks spanning medical decision support[39, 41], cross-modal inference[31, 5], temporal reasoning[40, 23], and natural language understanding[9, 21]. Sparse Mixture-of-Experts (SMoE) architectures[34, 35, 22] are widely used in multimodal learning for their ability to assign different experts to different input patterns through conditional computation. They are particularly effective when all input modalities are present, enabling diverse feature extraction and flexible expert routing. However, this assumption is often violated in practice. Incomplete modality observations are common in real-world scenarios due to sensor failures, privacy restrictions, or heterogeneous data collection pipelines. When one or more modalities are missing, SMoE models receive partial input signals, which undermines the reliability of the gating mechanism and results in suboptimal expert selection. Although some methods attempt to assign different modality combinations to different experts, this strategy becomes computationally intractable as the number of modalities grows exponentially. To reduce this complexity, recent SMoE frameworks typically activate the Top-K experts per sample. However, the use of softmax-based routing[9] tends to produce sharp distributions, often concentrating the routing on a single expert. This leads to the expert collapse problem [9, 12, 35], where only a few experts dominate the computation as shown in Figure 1(a), resulting in poor diversity and limited generalization. Efforts to mitigate this issue using load-balancing losses introduce additional optimization challenges. In particular, they may cause unstable or ambiguous expert selection during training, as illustrated in Figure3(a).
@@
While recent works[9, 12, 35] has tackled expert collapse within SMoE architectures, handling missing modalities remains an orthogonal and unresolved challenge. In the context of missing modalities, many recent approaches aim to reconstruct absent views through cross-modal information sharing [20, 29, 34] as shown in Figure 1(b). However, these methods typically introduce multiple auxiliary losses to jointly learn task-relevant, shared, and modality-specific features, which increases training complexity and may inadvertently erode modality-unique information. To alleviate this issue, recent work has proposed using learnable modality banks[35] to impute missing inputs based on the available modality combinations. Nevertheless, such banks are usually initialized randomly, similar to prompt tuning[15, 18], and often fail to capture fine-grained modality-specific structure, limiting their effectiveness in practical settings.
@@
To overcome the limited robustness of existing SMoE models under incomplete modality conditions, we propose Confidence-guided Sparse Mixture-Of-E (Conf-SMoE), a unified framework that introduces two novel mechanisms for handling arbitrary missing modalities. Our method stems from a core insight: modality-specific structure should be preserved and leveraged for both reconstruction and routing, a property often overlooked in prior work. Based on this, we propose a two-stage imputation strategy that first estimates missing modalities using intra-modality correlations, and then refines these estimates through instance-level cross-modal alignment, as illustrated in Figure 1(c). The rationale behind this approach is inspired by the modality gap phenomenon discussed by [19], which suggests that a certain level of discrepancy between modalities is both expected and beneficial for downstream tasks. Thus, it is natural to first impute missing modalities using intra-modality information subsequently enhance them through instance-level cross-modal refinement.
@@
We further uncover and formalize a critical cause of expert collapse in SMoE: the gradient concentration induced by softmax gating, which suppresses routing diversity and limits generalization[9, 30, 35, 12, 6]. Through theoretical analysis, we show that conventional load-balancing losses exacerbate this issue by introducing gradient conflicts during optimization, reflecting by the "Sinoual Wave" pattern in expert selection plot shown in Figure 3(a). To resolve this, we design a confidence-guided expert routing mechanism that replaces softmax gating with token-level confidence scores. This formulation not only avoids the adverse effects of entropy-based balancing but also enables stable, interpretable expert selection. By decoupling routing from the optimization bottlenecks of prior designs, Conf-SMoE achieves both enhanced expert diversity and consistent performance across a wide range of incomplete modality scenarios. Our contributions are the following:
@@ 2.1 Mixture-of-Expert Architecture
In this work, we build upon the SMoE framework by incorporating expert routing into the Transformer’s feedforward layers, enabling dynamic token-to-expert assignment based on learned routing scores. This architecture supports selective expert activation, where only the Top-K experts are engaged per token, and proves especially effective in the joint expert setting[12], allowing tokens from different modalities to be routed to shared experts for enhanced cross-modal interaction. We begin with the standard Softmax SMoEs as a baseline configuration. Denote that input token embedding $\mathbf{h}\in\mathbb{R}^{d}$ and the a learnable softmax router $G(\mathbf{h})=softmax(\mathbf{h})$ that produces routing score $g_{i}$ for each expert $E_{i}$. Then, Top-K function is applied to select the top-k expert in forward, the final tokens/samples representation is weighted summation from Top-K expert with corresponding routing score $g_{i}$. Formally, we can represent the forward process of SMoE as:
$E(\mathbf{h})=\mathbf{h}+\sum_{i=1}^{K}G(\mathbf{h})E_{i}(\mathbf{h})$ (1)
Where $E_{i}(\cdot)$ represent the i-th expert chosen by the router in expert pool and we would like to sum up the representation from Top-K expert by gating score out of $G(\mathbf{h})$.
@@ 2.2 What Causes Expert Collapse?
The Jacobian matrix of Eq. 1 with respect to $\mathbf{h}$ can be analyzed to reveal the underlying cause of expert collapse in SMoE. Therefore,
$\mathbf{J}_{MoE}=\sum_{i=1}^{K}\frac{\partial softmax(\mathbf{h})}{\partial\mathbf{h}}E_{i}(\mathbf{h})+\sum_{i=1}^{K}G(\mathbf{h})E_{i}^{{}^{\prime}}(\mathbf{h})$ (2)
Let’s denote $\mathbf{g}=softmax(\mathbf{u})$, where $\mathbf{u}\in\mathbb{R}^{N}$ is the gating logits to select expert. The Jacobian matrix for the softmax function is denoted as $\mathbf{J}_{softmax}\in\mathbb{R}^{d\times d}$, where each element $\mathbf{J}_{ij}=\frac{\partial softmax(\mathbf{h_{j})}}{\partial\mathbf{h_{j}}}=g_{i}(\delta_{ij}-g_{i})$ and $\delta_{ij}$ is Kronecker delta. In matrix form, we can express Jacobian of softmax as: $\mathbf{J}_{softmax}=\textbf{diag(g)}-\textbf{g}\textbf{g}^{T}$, and the set of expert forward function $\mathbf{E}_{K}(\mathbf{h})=[E_{1}(\mathbf{h}),E_{2}(\mathbf{h}),\dots,E_{K}(\mathbf{h})]\in\mathbb{R}^{d\times d}$. For simplicity, we simplify the Top-K operation into $\mathbf{E_{K}}$. Therefore, we can rewrite Eq. 2 to Eq. 3: the first term is to learn a better routing score driven by gradient of gating score $\mathbf{g}$ and second term is to learn a better representation as the gradient driven by $E_{i}^{{}^{\prime}}(\mathbf{h})$. The Top-k mechanism makes the first term of Eq. 3 sparse and only a few experts contribute to gradient while only one expert dominate the gradient due to softmax sharp distribution. It is clear that the greatest gating score from the sharp distribution of softmax dominates the gradient in training phrase, resulting expert collapse based on token preference and produce rich-get-richer expert.
$\mathbf{J}_{MoE}=\underbrace{\mathbf{E}_{K}(\mathbf{h})(\textbf{ diag(g) }-\textbf{g}\textbf{g}^{\top})}_{\text{Learning a better routing score}}\;\;\;\;\;+\underbrace{\sum_{i=1}^{K}G(\mathbf{h})E_{i}^{{}^{\prime}}(\mathbf{h})}_{\text{Learning a better representation}}$ (3)
@@ 2.3 Gradient of Load Balance Loss
This expert collapse phenomenon restrict the model learning capacity since many other expert are wasted and cannot reflect the actual expert diversity. Two widely-used solutions to achieve balance expert usage is to introduce additional load balance loss and expert capacity. However, such load balance loss approach is difficult to optimize since it produces gradient direction opposite to softmax routing score terms in Eq 3. We consider that any load balance loss $\mathcal{L}_{load}$ will eventually results a flat and diverse softmax distribution, which encourage higher entropy on the softmax distribution. From Assumption 1 defined in Appendix, we define that the objective of any $\mathcal{L}_{load}$ as $\frac{1}{\mathcal{H(\mathbf{g})}}$ over the softmax distribution, where $\mathcal{H(\mathbf{g})}=-\sum_{i=1}^{N}g_{i}log\ g_{i}$ represent the entropy of softmax distribution g. Taking the Jacobian over the $\mathcal{L}_{load}$ will give us (See detailed derivation in Appendix. B.2):
@@
Recall that $\mathbf{E}_{K}(\mathbf{h})$ in Eq. 3 is a set of expert with ReLU activation, we assume this network should not output negative value which is theoretically true in most cases. We know the fact that the Jacobian of softmax $\mathbf{J}_{softmax}$ is symmetry and always positive semi-definite[10], making the optimization of better routing score is always positive. In contrast, $log(\textbf{g})+1=[log(g_{1})+1,log(g_{2})+1,\dots,log(g_{K})+1]$. When sharp distribution appears, only one $g_{j}$ is closed to 1 while other $g_{i},(i\neq j)$ are closed to 0, which is practically and mostly true. Therefore, $\frac{1}{\mathcal{H(\textbf{g})}^{2}}(log(\textbf{g})+1)^{\top}$ is exponentially dominated by negative value and $\mathbf{J}_{load}$ is opposite to the first term of $\mathbf{J}_{MoE}$. Therefore, the opposite gradient from axillary load balance loss makes the training difficult to converge to global minimum and Expert selection becomes ambiguous, failing to reflect actual capacity of expert.
@@ 3.1 ConfNet:Confidence-Guided Gating Network
From Eq. 4 and Eq. 3, we understand that load balance loss shows opposite gradient direction to softmax router. Then, the potential solution to explicitly achieve load balance is to remove the association of sharp distribution from Eq. 3. One is to replace the routing function as illustrated in Figure. 2 (c) and one is to decouple the gating score with softmax router as illustrated in Figure. 2 (d). A typical example from a recent work, FuseMoE [12], they replace the Softmax router to Laplacian router and calculate the routing score based on similarity of the expert embedding $W_{N}$ by $\mathcal{L}_{1}$ distance. Then the Top-K experts are selected with highest similarity to expert embedding. The second approach is to decouple the connection of gating score and expert selection. One simple approach is to consider all expert are equally important, then the final representation is averaged from all selected expert. However, this averaged approach does not emphasize the difference of expert and lose the specification of expert.
@@
Therefore, we propose a Confidence-Guided Gating Network Pool, which introduces a novel routing strategy guided by task confidence. Specifically, instead of solely relying on conventional softmax-based gating, which often results in sharp distributions and entangled gradient contributions, our method leverages the ground-truth label information to supervise the routing mechanism, encouraging expert selection aligned with task-relevant confidence rather than token or embedding similarity. For each expert $E_{i}$, we introduce an auxiliary confidence estimation network: $U_{i}$ that maps token embedding $\mathbf{h}\in\mathbb{R}^{d}$ to a single scalar logits $v_{i}\in\mathbb{R}$. $v_{i}$ is then mapped to a probability $g_{i}\in(0,1)$ as gating score by Sigmoid, representing the confidence of ground truth label $y_{t}$. To learn the true confidence $p_{t}$ of downstream task w.r.t $y_{t}$, where $t$ is the notation for ground truth. We apply MSE loss as an axillary loss that minimize the difference between $g_{i}$ and $p_{t}$. Formally, given any classification dataset $D$ with supervised signal $y_{i}$ and number of instance $|D|$, we consider the task loss $\mathcal{L}_{task}$ as Cross Entropy Loss and the final objective is to optimize:
$\mathcal{L}=\mathcal{L}_{task}+\mathcal{L}_{conf}=\frac{1}{|D|}\sum^{D}p_{t}log(p_{t})+\frac{1}{|D|K}\sum^{D}\sum^{K}_{i=1}(g_{i}-p_{t})^{2}$ (5)
@@
The intuition is that the learning of gating score in Eq. 3 will be driven by the task-direct score, this forces model to select expert with task-preference instead of token preference as Softmax gating [9] or expert embedding preference as Laplacian gating does [12]. During inference, we use the learned confidence score $g_{i}$ from the ConfNet as a confidence proxy for the task-specific relevance of each expert. These scores are used for specialized token fusion, enabling dynamic integration of modality-specific and task-relevant information. The $g_{i}$ is Token-level confidence, which each token will have one corresponding $g_{i}$ for aggregation. Additionally, to capture coarse-grained signals, we aggregate token-level representations assigned to each expert and average them to compute expert-level gating signals as Expert-level variant, denoted as ConfSMoE-E in experiment. This confidence-driven routing enables both fine-grained specialization and robust multimodal interaction, leading to improved generalization and interpretability.
@@
Figure 2: The Framework of Conf-SMoE: (a) Conf-SMoE consider the confidence learned from ConfNet pool as the gating score for expert aggregation. Missing modality imputation is achieved by taking the mean of random instances from the corresponding modality of training data as common modality specific feature. The instance-specific feature are sparsely aggregated from existing modality. (b) Expert collapse is because of the softmax sharp gating score. (c)(d) "Fair" gating score distribution and decoupling gating score can alleviate the expert collapse without the helps of load balance loss.

### v2

@@ Abstract
Effectively managing missing modalities is a fundamental challenge in real-world multimodal learning scenarios, where data incompleteness often results from systematic collection errors or sensor failures. Sparse Mixture-of-Experts (SMoE) architectures have the potential to naturally handle multimodal data, with individual experts specializing in different modalities. However, existing SMoE approach often lacks proper ability to handle missing modality, leading to performance degradation and poor generalization in real-world applications. We propose ConfSMoE to introduce a two-stage imputation module to handle the missing modality problem for the SMoE architecture by taking the opinion of experts and reveal the insight of expert collapse from theoretical analysis with strong empirical evidence. Inspired by our theoretical analysis, ConfSMoE propose a novel expert gating mechanism by detaching the softmax routing score to task confidence score w.r.t ground truth signal. This naturally relieves expert collapse without introducing additional load balance loss function. We show that the insights of expert collapse aligns with other gating mechanism such as Gaussian and Laplacian gate. The proposed method is evaluated on four different real world dataset with three distinct experiment settings to conduct comprehensive analysis of ConfSMoE on resistance to missing modality and the impacts of proposed gating mechanism. We release our official implementation in: https://github.com/IcurasLW/Official-Repository-of-ConfSMoE.git
@@ 1 Introduction
As modern applications increasingly involve data from multiple heterogeneous sources, multimodal learning has emerged as a critical paradigm to unlock complementary insights and enhance decision-making in complex tasks spanning medical decision supportZheng et al. (2024a, c), cross-modal inferenceWu et al. (2017); Cheng et al. (2024), temporal reasoningZheng et al. (2024b); Ni et al. (2024), and natural language understandingFedus et al. (2022); Masoudnia and Ebrahimpour (2014). Sparse Mixture-of-Experts (SMoE) architecturesYao et al. (2024); Yun et al. (2024); Mustafa et al. (2022) are widely used in multimodal learning for their ability to assign different experts to different input patterns through conditional computation. They are particularly effective when all input modalities are present, enabling diverse feature extraction and flexible expert routing. However, this assumption is often violated in practice. Incomplete modality observations are common in real-world scenarios due to sensor failures, privacy restrictions, or heterogeneous data collection pipelines. When one or more modalities are missing, SMoE models receive partial input signals, which undermines the reliability of the gating mechanism and results in suboptimal expert selection. Although some methods attempt to assign different modality combinations to different experts, this strategy becomes computationally intractable as the number of modalities grows exponentially. To reduce this complexity, recent SMoE frameworks typically activate the Top-K experts per sample. However, the use of softmax-based routingFedus et al. (2022) tends to produce sharp distributions, often concentrating the routing on a single expert. This leads to the expert collapse problem Fedus et al. (2022); Han et al. (2024); Yun et al. (2024), where only a few experts dominate the computation as shown in Figure 1(a), resulting in poor diversity and limited generalization. Efforts to mitigate this issue using load-balancing losses introduce additional optimization challenges. In particular, they may cause unstable or ambiguous expert selection during training, as illustrated in Figure3(a).
@@
While recent worksFedus et al. (2022); Han et al. (2024); Yun et al. (2024) has tackled expert collapse within SMoE architectures, handling missing modalities remains an orthogonal and unresolved challenge. In the context of missing modalities, many recent approaches aim to reconstruct absent views through cross-modal information sharing Ma et al. (2021); Wang et al. (2023); Yao et al. (2024) as shown in Figure 1(b). However, these methods typically introduce multiple auxiliary losses to jointly learn task-relevant, shared, and modality-specific features, which increases training complexity and may inadvertently erode modality-unique information. To alleviate this issue, recent work has proposed using learnable modality banksYun et al. (2024) to impute missing inputs based on the available modality combinations. Nevertheless, such banks are usually initialized randomly, similar to prompt tuningJia et al. (2022); Lester et al. (2021), and often fail to capture fine-grained modality-specific structure, limiting their effectiveness in practical settings.
@@
To overcome the limited robustness of existing SMoE models under incomplete modality conditions, we propose Confidence-guided Sparse Mixture-Of-E (ConfSMoE), a unified framework that introduces two novel mechanisms for handling arbitrary missing modalities. Our method stems from a core insight: modality-specific structure should be preserved and leveraged for both reconstruction and routing, a property often overlooked in prior work. Based on this, we propose a two-stage imputation strategy that first estimates missing modalities using intra-modality correlations, and then refines these estimates through instance-level cross-modal alignment, as illustrated in Figure 1(c). The rationale behind this approach is inspired by the modality gap phenomenon discussed by Liang et al. (2022), which suggests that a certain level of discrepancy between modalities is both expected and beneficial for downstream tasks. Thus, it is natural to first impute missing modalities using intra-modality information subsequently enhance them through instance-level cross-modal refinement.
@@
We further uncover and formalize a critical cause of expert collapse in SMoE: the gradient concentration induced by softmax gating, which suppresses routing diversity and limits generalizationFedus et al. (2022); Wang et al. (2024); Yun et al. (2024); Han et al. (2024); Chi et al. (2022). Through theoretical analysis, we show that conventional load-balancing losses exacerbate this issue by introducing gradient conflicts during optimization, reflecting by the "Sinoual Wave" pattern in expert selection plot shown in Figure 3(a). To resolve this, we design a confidence-guided expert routing mechanism that replaces softmax gating with token-level confidence scores. This formulation not only avoids the adverse effects of entropy-based balancing but also enables stable, interpretable expert selection. By decoupling routing from the optimization bottlenecks of prior designs, ConfSMoE achieves both enhanced expert diversity and consistent performance across a wide range of incomplete modality scenarios. Our contributions are the following:
@@ 2.1 Mixture-of-Expert Architecture
In this work, we build upon the SMoE framework by incorporating expert routing into the Transformer’s feedforward layers, enabling dynamic token-to-expert assignment based on learned routing scores. This architecture supports selective expert activation, where only the Top-K experts are engaged per token, and proves especially effective in the joint expert setting Han et al. (2024), allowing tokens from different modalities to be routed to shared experts for enhanced cross-modal interaction. We begin with the standard Softmax SMoEs with $N$ number of expert as a baseline configuration. Denote that input token embedding $\mathbf{h}\in\mathbb{R}^{d}$ and the a learnable softmax router $G(\mathbf{h})=softmax(\mathbf{h})$ that produces routing score $g_{i}$ for each expert $E_{i}$ from expert pool $\mathbf{E}=[E_{1},E_{2},\dots,E_{N}]\in\mathbb{R}^{d\times d}$. Then, Top-K function is applied to select the Top-K experts in forward, the final tokens/samples representation is weighted summation from Top-K experts with corresponding routing score $g_{i}$. Formally, we can represent the forward process of SMoE as:
$E(\mathbf{h})=\mathbf{h}+\sum_{i}^{K}g_{i}E_{i}(\mathbf{h})$ (1)
@@ 2.2 What Causes Expert Collapse?
The Jacobian matrix of Eq. 1 w.r.t $\mathbf{h}$ can be analyzed to reveal the underlying cause of expert collapse in SMoE. Therefore,
$\mathbf{J}_{MoE}=\sum_{i}^{K}\frac{\partial g_{i}}{\partial\mathbf{h}}E_{i}(\mathbf{h})+\sum_{i}^{K}g_{i}E_{i}^{{}^{\prime}}(\mathbf{h})$ (2)
[…]
$\mathbf{J}_{MoE}=\underbrace{\mathbf{E}_{K}(\mathbf{h})(\textbf{diag(g)}_{K}-\textbf{g}\textbf{g}_{K}^{\top})}_{\text{Learning a better routing score}}\;\;\;\;\;+\underbrace{\sum_{i}^{K}g_{i}E_{i}^{{}^{\prime}}(\mathbf{h})}_{\text{Learning a better representation}}$ (3)
@@ 2.3 Gradient of Load Balance Loss
This expert collapse phenomenon restrict the model learning capacity since many other expert are wasted and cannot reflect the actual expert diversity. One widely-used solutions to achieve balance expert usage is to introduce additional load balance loss Yun et al. (2024); Han et al. (2024). However, such load balance loss approach is difficult to optimize since it produces gradient direction opposite to routing score terms in Eq 3. To show this, we consider that any load balance loss $\mathcal{L}_{load}$ will eventually results a diverse softmax routing distribution, which encourage higher entropy on the routing score distribution. From Assumption 1 defined in Appendix, we define that the objective of any $\mathcal{L}_{load}$ as $\frac{1}{\mathcal{H(\mathbf{g})}}$ over the softmax distribution, where $\mathcal{H(\mathbf{g})}=-\sum_{i=1}^{N}g_{i}log\ g_{i}$ represent the entropy of softmax distribution g. Taking the Jacobian over the $\mathcal{L}_{load}$ will give us (See detailed derivation in Appendix. B.2):
@@
Recall that $\mathbf{E}_{K}(\mathbf{h})$ in Eq. 3 is a set of expert with ReLU activation, we assume this network should not output negative value as activation is located on positive side. We also know the fact that the Jacobian of softmax $\mathbf{J}_{softmax}$ is symmetry and always Positive Semi-Definite (PSD) Gao and Pavel (2017), making the optimization of better routing score is always positive. In contrast, $log(\textbf{g})+1=[log(g_{1})+1,log(g_{2})+1,\dots,log(g_{N})+1]$. When sharp distribution appears, only one $g_{j}$ is closed to 1 while other $g_{i},(i\neq j)$, are closed to 0. Therefore, $\frac{1}{\mathcal{H(\textbf{g})}^{2}}(log(\textbf{g})+1)^{\top}$ is exponentially dominated by negative value, producing conflict gradient between $\mathbf{J}_{load}$ and the first term of $\mathbf{J}_{MoE}$. The Top-K operation does not change PSD property of term, but the gradient will be backward through the off Top-K term in $J_{MoE}$. It is also worthy to notice that gradient in Eq. 4 conflict occurs for all Top-K number as the load balance loss consider all $g_{i}$ regardless of number K expert selection. Taking an extreme example, Top-1 selection Fedus et al. (2022) would generate one $g_{i}$ out of g for the second term of Eq. 3 while Eq. 4 still take all $g_{i}$ to gradient computation. Therefore, the opposite gradient from axillary load balance loss makes the training difficult to converge to global minimum and Expert selection becomes ambiguous, failing to reflect actual capacity of expert.
@@ 3.1 ConfNet:Confidence-Guided Gating Network
From Eq. 4 and Eq. 3, we understand that load balance loss shows opposite gradient direction to softmax router in typical SMoE. To avoid this gradient conflict and balance the load, the potential solution is to remove the association of sharp distribution from Eq. 3 and attempt to explicitly achieve load balance. One is to replace the routing function as illustrated in Figure. 2 (c) and one is to decouple the gating score with softmax router as illustrated in Figure. 2 (d). A typical example from a recent work, FuseMoE Han et al. (2024), they replace the Softmax router to Laplacian router and calculate the routing score based on similarity of the expert embedding by $\mathcal{L}_{1}$ distance. Then the Top-K experts are selected with highest similarity to expert embedding and Laplacian router is not a sharp distribution so that the load is explicit balanced. The second approach is to decouple the connection of gating score and expert selection. One simple approach is to consider all expert are equally important, then the final representation is averaged from all selected expert. However, this averaged approach does not emphasize the difference of expert and lose the specification of expert.
@@
Therefore, we propose a Confidence-Guided Gating Network Pool, which introduces a novel routing strategy guided by task confidence. Specifically, instead of solely relying on conventional softmax-based gating, which often results in sharp distributions and entangled gradient contributions, our method leverages the ground-truth label information to supervise the routing mechanism, encouraging expert selection aligned with task-relevant confidence rather than token or embedding similarity as in Han et al. (2024); Fedus et al. (2022). For each expert $E_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$, we introduce an auxiliary confidence estimation network: a one-layer linear network $U_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{1}$ that maps token embedding $\mathbf{h}\in\mathbb{R}^{d}$ to a single scalar logits $v_{i}\in\mathbb{R}^{1}$. $v_{i}$ is then mapped to a confidence score $c_{i}\in(0,1)$ as gating score by Sigmoid, representing the confidence of ground truth label $y_{t}$. To learn the true confidence $p_{t}$ of downstream task w.r.t $y_{t}$, where $t$ is the notation for ground truth. We apply MSE loss as an axillary loss that minimize the difference between $c_{i}$ and $p_{t}$. Formally, given any classification dataset $D$ with supervised signal $y_{i}$ and number of instance $|D|$, we consider the task loss $\mathcal{L}_{task}$ as Cross Entropy Loss and the final objective is to optimize:
$\mathcal{L}=\mathcal{L}_{task}+\mathcal{L}_{conf}=\frac{1}{|D|}\sum^{D}y_{t}log(p_{t})+\frac{1}{|D|K}\sum^{D}\sum^{K}_{i=1}(c_{i}-p_{t})^{2}$ (5)
@@
Note that Eq. 5 is convex as both cross-entropy and MSE loss are well-known convex. During both training and inference, $c_{i}$ serves as a proxy gating score for each expert to token fusion in Eq. 2. The $c_{i}$ is Token-level confidence, which each token has one corresponding $c_{i}$ for aggregation as in $\sum^{K}_{i}G(\mathbf{h})E_{i}(\mathbf{h})$ as in Eq. 1, resulting in a variant of proposed method for fine-grained feature learning, denoted as ConfSMoE-T. Additionally, a expert-level gating signal is also implemented for coarse-grained information retrieval. We aggregate token-level representations assigned to each expert and average them to compute expert-level gating signals as Expert-level variant, denoted as ConfSMoE-E in experiment. The multimodal interaction is learned during the token assignment for all experts, which all modalities are shared one identical expert pool. This confidence-driven routing enables both fine-grained specialization and robust multimodal interaction, leading to improved generalization and interpretability.
@@
Figure 2: The Framework of ConfSMoE: (a) ConfSMoE consider the confidence learned from ConfNet pool as the gating score for expert aggregation. The ConfNet pool is a set of one-layer linear weight $\mathbf{U}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{1}$. Missing modality imputation is achieved by taking the mean of random instances from the corresponding modality of training data as common modality specific feature. The instance-specific feature are sparsely aggregated from existing modality. (b) Sharp gating score causes expert collapse. (c)(d) "Fair" gating score distribution and decoupling gating score can alleviate the expert collapse without the helps of load balance loss.